OpenStudy (anonymous):

Improper Integral 0∫-inf(1/(2x-1)^2))

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits_{-\infty}^0 \frac{1}{(2x-1)^2}dx\]Mmmm ok what part you having trouble with?

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits_{-\infty}^0 \frac{1}{(2x-1)^2}dx\quad = \quad \int\limits\limits_{-\infty}^0 (2x-1)^{-2}dx\]

zepdrix (zepdrix):

Power Rule for Integration. If the 2x is confusing you, you can do a `u-sub` I guess.

OpenStudy (anonymous):

Thanks zepdrix